### v2

# Addressing Hallucinations with RAG and NMISS in Italian Healthcare LLM Chatbots

Abstract
I combine detection and mitigation techniques to addresses hallucinations in Large Language Models (LLMs). Mitigation is achieved in a question-answering Retrieval-Augmented Generation (RAG) framework while detection is obtained by introducing the Negative Missing Information Scoring System (NMISS), which accounts for contextual relevance in responses. While RAG mitigates hallucinations by grounding answers in external data, NMISS refines the evaluation by identifying cases where traditional metrics incorrectly flag contextually accurate responses as hallucinations. I use Italian health news articles as context to evaluate LLM performance. Results show that Gemma2 and GPT-4 outperform the other models, with GPT-4 producing answers closely aligned with reference responses. Mid-tier models, such as Llama2, Llama3, and Mistral benefit significantly from NMISS, highlighting their ability to provide richer contextual information. This combined approach offers new insights into the reduction and more accurate assessment of hallucinations in LLMs, with applications in real-world healthcare tasks and other domains.
Keywords: Large Language Models, Retrieval-Augmented Generation, Hallucinations, Healthcare, Natural Language Processing, Evaluation Metrics, Information Systems
1 Introduction
Large Language Models (LLMs) are AI systems typically built on the Transformer architecture, as introduced by Vaswani [52]. Designed to understand and generate human language, LLMs are trained on extensive text datasets, allowing them to grasp the nuances of language. These models perform various tasks, from basic text classification to advanced text generation, and more recent iterations can even interpret images and generate code.
LLMs have surged in popularity due to their intuitive, prompt-based interfaces, enabling effortless interaction. The past few years have seen the continuous development of notable LLMs like ChatGPT [33, 1], Gemini [11], Falcon [37], Llama [49, 50, 30], and Gemma [46, 47]. As highlighted by [15], these models excel in language comprehension [14, 16], text summarization [59], and reasoning [54, 22, 38]. As a result, their applications span various industries, such as processing electronic medical records, matching clinical trials, and drug discovery in healthcare. In finance, LLMs assist with fraud detection and trading, while in business, they automate customer service via chatbots and virtual assistants [35].
However, LLMs face a significant challenge: hallucinations [12, 51]. Hallucinations occur when models generate nonsensical or inaccurate content, failing to reflect the source material accurately [17]. This behavior is a serious threat since it affects the application of LLMs to real-world cases [20], as hallucinations can undermine the reliability of their outputs. [15] classifies hallucinations into two types: factuality hallucinations, where generated content diverges from verifiable facts, and faithfulness hallucinations, where the content strays from user instructions or input context. Factuality hallucinations manifest as either factual inconsistencies or fabrications. For example, a model might mistakenly claim that “The Colosseum is located in Berlin, Germany” instead of Rome or invent details about non-existent technologies. Faithfulness hallucinations occur when the output deviates from user instructions or context. These can involve instruction inconsistencies (e.g., responding in Spanish when asked to translate into Italian), context inconsistencies (introducing unrelated information), or logical errors (e.g., miscalculating “two times four” as three).
The mitigation of hallucinations has become an urgent concern due to the widespread use of LLMs in sensitive areas such as healthcare. Addressing hallucinations is challenging due to LLMs’ reliance on reinforcement learning from human feedback [34], their opaque black-box nature that hides knowledge boundaries [45, 42], and their training on large datasets that may contain outdated or incorrect information [60]. Additionally, their versatility across many tasks complicates mitigation efforts, while their strong performance can make hallucinations, especially factual fabrications, harder to detect.
Researchers focused on various techniques to mitigate hallucinations. One of the most promising methods is Retrieval-Augmented Generation (RAG) [24], which incorporates external knowledge during text generation to anchor outputs in factual information. RAG addresses issues of accuracy and currency in LLM responses by using authoritative sources, generating verifiable answers, and enhancing user confidence across applications [21]. Other approaches include Prompt Engineering that involves optimizing LLMs outputs with different instructions [48, 55], self-refinement through feedback and reasoning [27], where feedback on LLM outputs improves the accuracy and quality of subsequent responses, and prompt tuning, which modifies instructions during fine-tuning to improve task-specific performance [23].
Mitigating hallucinations requires first their detection, and various metrics have emerged. Traditional metrics like ROUGE [26], BLEU [36], and METEOR [4] measure consistency between generated text and source text, but they often struggle with semantically relevant but superficially divergent text [5]. To address these limitations, newer embedding-based methods such as BERTScore [58] use contextual embeddings to assess semantic similarity between sentences. Another approach involves ”opening the box” where token probabilities or logit outputs of LLMs are analyzed to detect hallucinations [3, 51]. However, these techniques require access to internal model states, which is not always feasible for proprietary systems.
This paper proposes a general framework for detecting and mitigating hallucinations, with an application to the Italian healthcare domain. For the mitigation part, I employ a RAG architecture to integrate external knowledge into the generation process, ensuring that the models are grounded in factual data. The methodology applies a multilingual Sentence Transformer model [40, 41] for vector database generation and retrieval. I compare a range of LLMs, including Gemma [46], Gemma 2 [47], Mistral [18], GPT-4 [1], Llama2 [50], and Llama3 [30] for the answer generation. This framework uses a prompting strategy to instruct models to generate answers based on the provided context. To assess interpretability and difficulty, I categorize 100 questions into three levels of understanding: low, mid, and high. For the detection strategy, I introduce the Negative Missing Information Scoring System (NMISS), a novel scoring framework designed to be applied alongside existing metrics to integrate contextual relevance. By analyzing residual words from machine-generated answers after comparison with human responses, NMISS refines traditional metrics by highlighting cases where they may unfairly penalize performance despite the enhanced contextual relevance of the responses.
Results show that Gemma2 and GPT-4 achieve the highest overall performance in Italian healthcare-related question-answering tasks across all difficulty levels, while Gemma exhibits the highest rate of hallucinations. The NMISS scoring system provides a more accurate assessment of LLM performance, especially in high-difficulty tasks for mid-tier models like Llama2, Llama3, and Mistral, which show notable improvements compared to traditional metrics. Additionally, NMISS recognizes the value of contextual information provided by models in cases where they fail to deliver exact matches to expected answers in difficult tasks. By focusing on the relevance and coherence of the provided context, NMISS highlights its advantage in evaluating responses that are partially correct or contextually appropriate, even when they do not fully align with the reference answers.
This paper is organized as follows: Section 2 reviews the relevant literature on hallucination detection and mitigation. Section 3 details the methodology, highlighting the contributions of NMISS. Section 4 describes the experimental setup on Italian healthcare news data, and Section 5 presents my findings. Finally, Section 6 summarizes contributions and outlines future research directions.
2 Related Works
2.1 Hallucination Detection Techniques
Hallucinations in LLMs have been widely studied. [15] provides a detailed overview of strategies to detect and mitigate hallucinations, offering insight into their occurrence in current LLMs architectures. Ensuring LLM’s trustworthiness depends on early detection, as emphasized by [51], who highlights the importance of proactive measures during text generation.
[60] categorize detection methods into three main types. First, logit-based estimation, which uses token-level probabilities or entropy from the model’s logits [39]. Second, verbalized confidence estimation, where LLMs are asked to self-rate the confidence in their responses. Third, consistency-based estimation involves generating multiple candidate outputs and comparing them to detect logical inconsistencies, mimicking human reasoning to refine answers [53, 44, 61].
Furthermore, [19] categorize hallucination detection into white-box, grey-box, and black-box techniques. White-box methods analyze the model’s internal workings to identify when incorrect statements are generated. Grey-box approaches examine logit outputs to find areas of low confidence in the output. Black-box techniques, which only require the model’s final answers, are simpler to apply since they don’t require access to internal states.
Recent studies introduce SAPLMA [3], which detects hallucinations by using a classifier trained to predict the truthfulness of statements based on LLMs’ hidden layer activations. This approach outperforms traditional methods that rely solely on sentence probabilities. [51] propose an iterative process that detects potential hallucinations by identifying tokens with low normalized probabilities. The system then retrieves relevant information to correct these errors, effectively reducing hallucinations in GPT-3.5.
2.2 Resource-Intensive Mitigation Strategies
In addition to detection, various strategies have been developed to mitigate hallucinations in LLMs. One common approach is Knowledge Augmentation, such as Knowledge Injection (KI). [29] enhance prompts with information from knowledge graphs, such as adding details like location names for review tasks. This significantly reduces hallucinations and improves factual accuracy, with KI-LLMs producing more correct assertions and outperforming larger models in response quality.
[31] take a direct approach by training T5 models on knowledge graph triples from Wikidata. By masking key entities, the model learns and generalizes knowledge effectively. SKILL, after fine-tuning, outperforms standard T5 models on question-answering tasks, proving the effectiveness of structured data infusion.
Another key mitigation strategy is retrieval-based frameworks. [25] combines a sequence-to-sequence model with a dense vector index of Wikipedia articles using the RAG framework. This integration enhances the accuracy and reliability of generated content. Similarly, [13] apply retrieval techniques in biomedical contexts, incorporating sources like UMLS and Wikipedia to generate more factual summaries, which are crucial for healthcare applications.
[62] introduce a self-supervision framework that assigns risk scores, called POLAR scores, to LLM responses as indicators of hallucination risk. Their approach uses Pareto optimization to balance supervision from multiple sources. The process involves gathering LLM responses, training a probabilistic function $h$ to estimate error rates, and optionally re-prompting the LLM based on the POLAR score with RAG support. Experimental results show significantly reduced hallucination rates across tasks like question answering and text generation.
[57] propose the MixAlign framework, which aligns user queries with external knowledge to resolve ambiguities. This improves response accuracy and reduces hallucinations, with results showing a sharp increase in correct answers and a decrease in hallucinations.
2.3 Zero-Resource Hallucination Strategies
The strategies discussed in Section 2.2 are resource-intensive. In contrast, zero-resource hallucination methods are designed for black-box models and scenarios with limited resources, relying only on model outputs without needing internal access or external data. Many of these techniques not only detect hallucinations but also help mitigate them by refining responses based on internal consistency assessments.
One approach is through self-consistency techniques, which compare multiple responses generated by the LLM itself. [28] propose a zero-resource method that generates multiple responses from GPT-3 and compares them to a reference answer to assess factual consistency. This method detects and mitigate hallucinations using low-level metrics and outperforms traditional grey-box and black-box methods.
Another approach is self-rating techniques, as proposed by [6]. The authors introduce RelD, a discriminator designed to detect hallucinations in LLM-generated answers. RelD is trained by converting a regression task into a binary classification task, aligning with human intuition. By flagging low-confidence responses, it serves as an indirect mitigation mechanism. Experiments across various datasets and models (e.g., BLOOM [56]) show RelD’s effectiveness, validated by both automatic metrics and human-in-the-loop evaluations. Similarly, [2] addresses hallucinations in fabricated references using Direct Queries (DQs) and Indirect Queries (IQs) without external resources. DQs involve asking the model a binary question multiple times to calculate a probability. In contrast, IQs ask about the authors or content of the reference and check for consistency across responses. Although primarily designed for detection, these methods can also reduce hallucinations by discarding low-confidence responses. Results show that IQs generally perform better than DQs in detecting hallucinations, though performance varies among models like GPT-4, GPT-3, and Llama 2.
Finally, [32] explore self-contradiction techniques as a factor in hallucinations. They introduce a prompting framework to detect and mitigate contradictions within the same response, relying solely on the model’s logical reasoning capabilities. This framework includes functions for gathering contextual information, generating sentences, identifying contradictions, and refining sentences. Across various LLMs, this approach achieves an 80% F1 score.
2.4 Metrics for Evaluating Hallucinations
Traditional metrics to detect or manage hallucinations [5] are notably ROUGE [26], BLEU [36], and METEOR [4]. These indicators measure the consistency between the generated text and the source text. However, a drawback of these simple metrics is that they do not accurately capture distance word-dependencies or if the two answers diverge in lexicon but have equal meaning. BERTScore overcomes this issue [58] by leveraging pre-trained contextual embeddings from BERT [9] and matches words in candidate and reference sentences by cosine similarity, correlating well with human judgments.
2.5 Contributions
Despite significant progress in hallucination detection and mitigation, key challenges persist. Many existing methods, such as knowledge injection or retrieval, are resource-intensive, limiting their practicality in certain applications. Additionally, traditional evaluation metrics often fail to account for contextual accuracy, unfairly penalizing semantically correct responses that differ lexically from reference answers.
This paper addresses these gaps in several steps. The key contributions of this work are as follows:
(1) A combined approach for hallucination detection and mitigation in LLMs. I develop a RAG-based architecture to integrate external knowledge into the generation process, ensuring that models are grounded in factual data. The methodology leverages a multilingual Sentence Transformer model [40, 41] for vector database generation and retrieval. This framework includes a prompting strategy to instruct models to generate answers based on the provided context, enabling a comparative evaluation of various LLMs using this approach.
(2) Comprehensive analysis of hallucinations across reasoning levels. To evaluate hallucination tendencies in LLMs, I analyze their performance across progressively challenging levels of reasoning, categorized in low, mid and high levels. This approach enables an assessment of how hallucination rates and mitigation effectiveness vary as tasks demand increasing degrees of logical complexity and understanding. The progressive evaluation highlights how model reliability and factual grounding are affected under escalating reasoning challenges, providing insights into their strengths and vulnerabilities.
(3) Development of a detection scoring system. I introduce the Negative Missing Information Scoring System (NMISS), a novel framework that fine-grains traditional metrics by incorporating contextual relevance. By analyzing residual words from machine-generated answers after comparison with human responses, NMISS highlights cases where traditional metrics may penalize models despite the enhanced contextual relevance of their answers.
3 Methodology
3.1 RAG Architecture
RAG models combine the strengths of retrieval-based and generative language models to address issues like hallucinations and factual inaccuracies by integrating a retrieval mechanism to ground responses in relevant information [24].
I formally define the problem as follows. Let $q$ represent the query, $c$ the context information retrieved from external sources, and $r$ the output sequence generated by the model. The architecture consists of two components: (i) a retriever $p_{\eta}(c|q)$ with parameters $\eta$, which selects a top-K set of documents $k$ based on the query $q$, and (ii) a generator $p_{\theta}(r_{i}|q,c,r_{1:i-1})$, with parameters $\theta$, which generates each token $r_{i}$ conditioned on the query $q$, retrieved context $c$, and previously generated tokens $r_{1:i-1}$.
$p_{\text{RAG}}(r\mid q)=\sum_{c\in k(p(\cdot\mid q))}p_{\eta}(c\mid q)\prod_{i=1}^{N}p_{\theta}\left(r_{i}\mid q,c,r_{1:i-1}\right)$ (1)
This RAG-sequence approach ensures that the response remains consistent by using the same retrieved document to generate each token, maintaining coherence throughout the sequence. Here, $N$ represents the length of the output sequence. This architecture is particularly advantageous for domains like healthcare, where factual accuracy is critical. By leveraging retrieval from medical documents, guidelines, or research articles, the RAG model ensures that responses are accurate, minimizing the risks associated with misinformation [10]. In this paper, I employ the retrieval and generation models separately and combine them for inference; the workflow is detailed in Section 4.2.
3.2 Evaluation Metrics
The metrics described in this section are used to evaluate the quality of responses generated by the system, referred to as the System Summary (SS), against the ground truth responses, referred to as the Reference Summary (RS), which are human-generated responses in this study. Both SS and RS are treated as sets of n-grams for all metrics. Formally, let $S_{\text{SS}}=\{s_{1},s_{2},\dots,s_{l}\}$, where $s_{i}$ is an n-gram in the SS, and $S_{\text{RS}}=\{r_{1},r_{2},\dots,r_{m}\}$, where $r_{j}$ is an n-gram in the RS.
ROUGE [26] evaluates responses based on precision, recall, and F1-score. The text is pre-processed before calculating these scores [43]. More formally, ROUGE-N measures the overlap of n-grams between the SS and RS. The precision, recall, and F1-score for ROUGE-N are defined as follows:
$\text{Precision}=\frac{|S_{\text{SS}}\cap S_{\text{RS}}|}{|S_{\text{SS}}|}$ (2)
$\text{Recall}=\frac{|S_{\text{SS}}\cap S_{\text{RS}}|}{|S_{\text{RS}}|}$ (3)
[…]
where $S_{\text{SS}}$ is the set of n-grams in the SS, $S_{\text{RS}}$ is the set of n-grams in the RS, $|S_{\text{SS}}\cap S_{\text{RS}}|$ is the cardinality of the intersection of $S_{\text{SS}}$ and $S_{\text{RS}}$, $|S_{\text{SS}}|$ is the cardinality of the set $S_{\text{SS}}$, and $|S_{\text{RS}}|$ is the cardinality of the set $S_{\text{RS}}$. Precision measures the accuracy of the generated text, while recall evaluates the coverage of the information. Low precision suggests hallucination, while low recall indicates a coverage issue.
Another variant used here is ROUGE-L, which measures the Longest Common Subsequence (LCS) between the SS and RS. The LCS represents the longest subsequence in both SS and RS while preserving word order, though the words do not need to appear consecutively [7].
BLEU [36] is designed to evaluate the quality of SS responses by comparing them against one or more RS answers. BLEU accounts for variability in word choice and word order found in different, yet valid, translations of the same source sentence. It quantifies the overlap by measuring how many n-grams from the SS are present in the RS. The core of the BLEU metric is the modified n-gram precision, which addresses the over-generation problem by treating each reference word as exhausted once a matching candidate word is found.
Let $V$ be a function that counts the occurrences of an n-gram in a given set of n-grams. Formally, for an n-gram $w$ and a set of n-grams $S$, the function $V$ is defined as $V(w,S)=|\{s\in S\mid s=w\}|$. This function returns the number of times $w$ appears in the set $S$. The modified n-gram precision $p_{n}$ is calculated as:
$p_{n}=\frac{\sum_{w\in S_{\text{SS}}}\min\left(V(w,S_{\text{SS}}),\max_{j}(V(w,S_{\text{RS}_{j}}))\right)}{\sum_{w\in S_{\text{SS}}}V(w,S_{\text{SS}})}$ (5)
where $V(w,S_{\text{SS}})$ is the number of times the n-gram $w$ appears in SS, and $\max_{j}(V(w,S_{\text{RS}_{j}}))$ is the maximum number of times $w$ appears across all RS. Here, $j$ iterates over the different reference summaries in $S_{\text{RS}}$, where $j=1,2,\dots,m$, and $m$ is the number of RS.
To account for translations that are too short, BLEU applies a brevity penalty (BP), defined as:
$\text{BP}=\begin{cases}1&\text{if }b>r\\ e^{(1-r/b)}&\text{if }b\leq r\end{cases}$ (6)
where $b$ is the length of the candidate translation (SS), and $r$ is the effective reference length (RS).
The final BLEU score is computed by taking the geometric mean of the modified n-gram precisions (up to length $N$, typically set to 4), and multiplying by the brevity penalty:
$\text{BLEU}=\text{BP}\cdot\exp\left(\frac{1}{N}\sum_{n=1}^{N}\log p_{n}\right)$ (7)
METEOR [4] addresses several limitations of BLEU by emphasizing word-to-word alignment between the SS and RS, accounting for exact matches, stems, and synonyms. METEOR computes the score by comparing the SS to each RS and reporting the highest score. Additionally, METEOR introduces a Discontiguous Matches Penalty (DP), based on the number of matched chunks.
Let $T$ represent the number of contiguous chunks of matched unigrams between SS and RS. For a set of matched unigrams $V=\{v_{1},v_{2},\dots,v_{k}\}$, a chunk is defined as a maximal sequence of consecutive matches $v_{i},v_{i+1},\dots,v_{i+t}$, such that $v_{i+t}=v_{i}+t$ for all $t\geq 0$.
The function $T(V,S_{\text{SS}})$, which counts the number of contiguous chunks, is then formalized as:
$T(V,S_{\text{SS}})=\left|\{\{v_{i},v_{i+1},\dots,v_{i+t}\}\subseteq V\mid v_{i+t}=v_{i}+t\}\right|$ (8)
This function returns the number of contiguous chunks of matched unigrams in $S_{\text{SS}}$. The Discontiguous Matches Penalty (DP) is then computed as:
$\text{DP}=0.5\cdot\left(\frac{T(V,S_{\text{SS}})}{|V|}\right)$ (9)
where $|V|$ is the number of matched unigrams. The penalty increases as the number of chunks grows, penalizing fragmented matches. The final METEOR score is adjusted by applying DP to the $F_{\text{mean}}$ score:
$\text{METEOR}=F_{\text{mean}}\cdot(1-\text{DP})$ (10)
BERTSCORE is an evaluation metric for text generation that leverages pre-trained BERT embeddings [58]. Unlike traditional metrics that rely on exact n-gram matches, BERTSCORE measures the semantic similarity between tokens in the System Summary (SS) and Reference Summary (RS) by comparing their contextual embeddings. This method captures the semantic meaning even when the exact words differ. BERTSCORE generates contextual embeddings for the tokens in both the SS and RS, where each embedding reflects the surrounding context of the word. The similarity between tokens is computed using cosine similarity. For each token in the RS, BERTSCORE identifies the most similar token in the SS, and vice versa, to compute precision and recall.
Given a RS sentence $x=\{x_{1},\ldots,x_{k}\}$ and a SS sentence $\hat{x}=\{\hat{x}_{1},\ldots,\hat{x}_{l}\}$, the token embeddings are represented as $\{e(x_{1}),\ldots,e(x_{k})\}$ and $\{e(\hat{x}_{1}),\ldots,e(\hat{x}_{l})\}$, where $e(\cdot)$ represents the contextual embedding for each token. The similarity between two embeddings is computed using cosine similarity. Precision ($P_{\text{BERT}}$) and recall ($R_{\text{BERT}}$) are calculated as follows:
$R_{\text{BERT}}=\frac{1}{|x|}\sum_{x_{i}\in x}\max_{\hat{x}_{j}\in\hat{x}}\cos(e(x_{i}),e(\hat{x}_{j}))$ (11)
$P_{\text{BERT}}=\frac{1}{|\hat{x}|}\sum_{\hat{x}_{j}\in\hat{x}}\max_{x_{i}\in x}\cos(e(\hat{x}_{j}),e(x_{i}))$ (12)
The final BERTSCORE is computed as the F1 score of the precision and recall:
$F_{\text{BERT}}=\frac{2\cdot P_{\text{BERT}}\cdot R_{\text{BERT}}}{P_{\text{BERT}}+R_{\text{BERT}}}$ (13)
Exact Match (EM) is a straightforward metric that penalizes any response not exactly equal to the provided dataset annotation. It is analogous to conventional classification accuracy, ensuring the generated text matches the reference text precisely. The EM can be defined as:
$EM=\begin{cases}1&\text{if }SS=RS\\ 0&\text{if }SS\neq RS\end{cases}$ (14)
3.3 NMISS
3.3.1 Rationale
The metrics discussed above evaluate the quality of text summaries and machine-generated responses by comparing them to reference answers. While effective, this method overlooks the possibility that non-matching words or n-grams might still be valuable if they align with the question’s broader context. This observation prompts an investigation into whether the words or phrases that do not match the reference could still be relevant compared to the broader context in which the response is generated. LLMs often produce responses beyond the strict instructions, frequently adding extra content, as shown in Table 1. This additional content can manifest in two ways: as hallucinations, where the model generates information that is factually inaccurate or irrelevant, or as meaningful elaborations, where the model extends the response in a contextually relevant manner.
Table 1: Example of Question, Context, Reference, and System Answer
Question How to prevent influenza?
Context [1] of the Relaunch Decree to: 1. prioritize investment in prevention to minimize impacts at all levels; 2. include in the Plans for managing a possible (re)emergence of an epidemic the entirety of health and social services to create the possibility of continuity of care even in the event of a “resurgence” of the epidemic (like a “generator group”), not only in hospitals but also in the community; 3. ensure that investments, particularly in personnel, cover all measures aimed at strengthening containment measures to reduce the risk of contagion. People should internalize and adopt recommendations such as washing hands thoroughly and frequently, using masks extensively, correctly, and appropriately, avoiding unnecessary movements and behaviors linked to social and recreational activities that could create spontaneous gatherings [3] as it is essential to maintain social distancing, use masks, and sanitize hands carefully to contain contagion. It is also crucial to self-isolate in case of cold symptoms. There are no risk locations but rather risky behaviors. We must convey to citizens the awareness that if the necessary precautions are followed, it is difficult to contract and transmit this infection, and most normal social and work activities can continue. [4] For the first time, high levels of criticality regarding the virus’s spread in our country are reported. The population is urged to avoid events and gatherings in both public and private settings. Individual strict behavior is mandatory to limit transmission risks and prevent a further and faster worsening of the epidemic. The local transmission of the virus is widespread across the entire territory.
Reference I’m sorry, I can’t help you based on the information I have.
System I’m sorry, I can’t help you based on the information I have, as the question is not related to the Relaunch Decree but rather to influenza, which is a different type of illness than the epidemic under discussion. However, based on the context, it is noted that the Relaunch Decree emphasizes the importance of strict behaviors to minimize impacts at all levels, such as social distancing, mask use, and hand hygiene, which are generally useful measures for preventing disease transmission in general, not just the discussed epidemic.
By accounting for these non-matching elements, I can more effectively evaluate the performance of LLMs, distinguishing between irrelevant or hallucinated information and valuable, contextually appropriate elaborations. This approach offers a more nuanced understanding of the model’s output.
The NMISS scoring system was developed to capture the potential value in these unmatched words by evaluating whether they align with the contextual information provided. This rationale stems from the observation that LLMs often do not strictly adhere to the given instructions, incorporating additional content into their answers. This content may either be a hallucination or a contextually relevant elaboration that enriches the response.
3.3.2 Formulation
Let =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ represent the set of reference answers, $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$ the set of system-generated answers, and $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$ the set of context texts associated with each reference answer. Define the set of matching tokens between system-generated answers and reference answers as $\mathcal{P}(s_{i},r_{i})$, and let $\mathcal{N}(s_{i},r_{i})=s_{i}\setminus\mathcal{P}(s_{i},r_{i})$ represent the tokens in $s_{i}$ that do not match $r_{i}$. Finally, let $\mathcal{A}(s_{i},c_{i})=\mathcal{N}(s_{i},r_{i})\cap c_{i}$, which represents the set of unmatched tokens in $s_{i}$ that match the context $c_{i}$. I aim to define a function
$f_{NMISS}:\mathcal{S}\times\mathcal{R}\times\mathcal{C}\rightarrow\mathbb{R}$ (15)
that evaluates how well system-generated answers align with both reference answers and the context. First, I define a reference-based evaluation function $f_{\text{ref},i}(s_{i},r_{i})=\text{X}_{\text{ref}}(\mathcal{P}(s_{i},r_{i}))$, and a context-based evaluation function $f_{\text{cxt},i}(s_{i},c_{i})=\text{X}_{\text{cxt}}(\mathcal{A}(s_{i},c_{i}))$, which computes the score based on tokens matching the context. The NMISS evaluation function is defined as:
$f_{\text{NMISS},i}(s_{i},r_{i},c_{i})=\max\left(f_{\text{ref},i}(s_{i},r_{i}),\frac{\lambda_{1}\cdot f_{\text{ref},i}(s_{i},r_{i})+\lambda_{2}\cdot f_{\text{cxt},i}(s_{i},c_{i})}{\lambda_{1}+\lambda_{2}}\right)$ (16)
where $\lambda_{1}=|\mathcal{P}(s_{i},r_{i})|$ and $\lambda_{2}=|\mathcal{A}(s_{i},c_{i})|$ represent the cardinalities of their respective sets. $\lambda_{1}$ and $\lambda_{2}$ represent the number of matching reference and context matches tokens, respectively.
To safeguard the integrity of standard scoring metrics, NMISS implements two key mechanisms: a weighted scheme and a maximum function. These mechanisms are essential for balancing reference-based and context-based contributions, addressing challenges such as score underestimation or overestimation when processing contextual information. The weighted scheme ensures a fair combination of reference-based and context-based scoring, preventing a disproportionate reliance on one over the other. Specifically, $\lambda_{1}$ and $\lambda_{2}$ control the relative influence of reference overlap and contextual alignment. This mechanism helps NMISS capture instances where a system-generated response extends beyond the reference but remains faithful to the context, distinguishing it from hallucinated content. The maximum function further enhances robustness by preventing NMISS from being unfairly penalized when context length inflates the denominator. In recall-sensitive metrics, an inflated denominator—such as the total number of n-grams in the context—could lower the score even if the additional content is meaningful. The formulation in Equation 16 ensures that context-based contributions only positively influence the score when they genuinely add value. In other words, if context matching introduces more noise than useful content, NMISS defaults to the reference-based score. This design prevents cases where large but irrelevant contexts artificially suppress the metric. This approach is particularly advantageous in tasks like question-answering, where the model’s response might not strictly match the reference but still be valid given the context. By integrating these mechanisms, NMISS ensures coherence in scoring, making it a more flexible and context-aware evaluation metric. The detailed implementation is outlined in Algorithm 1.
Algorithm 1 NMISS Algorithm
1:=$\mathcal{R}=\{r_{1},r_{2},\ldots,r_{n}\}$, the set of reference answers
2:$\mathcal{S}=\{s_{1},s_{2},\ldots,s_{n}\}$, the set of system-generated answers
3:$\mathcal{C}=\{c_{1},c_{2},\ldots,c_{n}\}$, the set of context texts
4:=$\mathcal{M}=\{m_{1},m_{2},\ldots,m_{n}\}$, the set of NMISS scores for a metric $\mathcal{X}$
5:
6:for each $(s_{i},r_{i},c_{i})$ in $(\mathcal{S},\mathcal{R},\mathcal{C})$ do
7:     Stage 1: Reference-Based Metric Calculation
8:     $\mathcal{P}(s_{i},r_{i})\leftarrow s_{i}\cap r_{i}$triangleright Matched tokens
9:     $\mathcal{N}(s_{i},r_{i})\leftarrow s_{i}\setminus\mathcal{P}(s_{i},r_{i})$triangleright Unmatched tokens
10:     $f_{\text{ref},i}(s_{i},r_{i})\leftarrow\text{X}_{\text{ref}}(\mathcal{P}(s_{i},r_{i}))$
11:
12:     Stage 2: Context-Based Metric Calculation
13:     $\mathcal{A}(s_{i},c_{i})\leftarrow\mathcal{N}(s_{i},r_{i})\cap c_{i}$triangleright Context-matched tokens
14:     $f_{\text{cxt},i}(s_{i},c_{i})\leftarrow\text{X}_{\text{cxt}}(\mathcal{A}(s_{i},c_{i}))$
15:
16:     Stage 3: NMISS Calculation
17:     $f_{\text{wgt},i}(s_{i},r_{i},c_{i})\leftarrow\frac{\lambda_{1}\cdot f_{\text{ref},i}(s_{i},r_{i})+\lambda_{2}\cdot f_{\text{cxt},i}(s_{i},c_{i})}{\lambda_{1}+\lambda_{2}}$
18:     $f_{\text{NMISS},i}(s_{i},r_{i},c_{i})\leftarrow\max(f_{\text{ref},i}(s_{i},r_{i}),f_{\text{wgt},i})$
19:     $\mathcal{M}_{\text{NMISS}}\leftarrow\mathcal{M}_{\text{NMISS}}\cup f_{\text{NMISS},i}(s_{i},r_{i},c_{i})$
20:end for
4 Experimental Setup
4.1 Data
Data comprises text-based news in Italian, collected from various sources, and all focused on the health domain within Italy. This collection encompasses regional updates, significant scientific developments, national headlines, and news on research and development, providing comprehensive coverage of health-related issues across the country. I collected 126,470 news articles from 2010 to 2024; their frequency is shown in Figure 1. A pre-processing step was employed to clean the raw text of HTML tags and other unwanted artifacts to ensure data quality. Token statistics are calculated to understand the overall structure of the data and are shown in Table 2. On average, each article contains approximately 348 tokens, with a standard deviation of 264 tokens, indicating some variability in article length. The shortest article consists of just 6 tokens, while the longest contains up to 5,594 tokens. The median article length is 284 tokens, with the interquartile range spanning from 195 to 421 tokens, reflecting a diverse range of article sizes.
Table 2: Summary of Token Statistics for the News Collection
N. Tokens Mean Std Min 25% Median 75% Max
44018696 348 264 6 195 284 421 5594
Figure 1: News Frequency Aggregated by Year
Next, I developed a list of 100 questions related to public health. Questions range from very generic to specific or ambiguous, leading to different levels of interpretation ease. The questions are categorized into three levels of understanding, based on the objective criterion of question complexity, measured by the average question length as a proxy for how detailed the question is (low: 30 characters, mid: 40 characters, high: 50 characters). Additionally, all questions were reviewed by a domain expert to ensure accuracy and relevance. All questions have been answered manually; Table 3 shows an example of questions.
Table 3: Categorized Example Questions
Category Example Question
HIGH How can epigenetics help us understand diseases?
MID What happens to the body when we are stressed?
LOW What is a common cold?
4.2 User-Agent Chatbot Framework
This section outlines the architecture of the user-agent chatbot designed to handle queries related to Italian health news, using a RAG approach that integrates a Dense Vector Store (DVS) with LLMs fine-tuned for Italian language processing. This multi-model setup ensures robustness in handling diverse queries related to Italian health news, providing accurate, coherent, and contextually grounded responses.
The system begins with Sentence Transformers to generate dense vector embeddings for the Document Database, which are stored and accessed via the Dense Vector Space. This enables efficient retrieval of relevant information through semantic search and clustering. After experimenting with different chunk lengths, I optimized retrieval performance by selecting a chunk length of 512 tokens with a 64-token overlap, as evaluated by human reviewers on a dataset of 100 health-related queries. For the retrieval part, I choose the Sentence Transformer model ‘paraphrase-multilingual-MiniLM-L12-v2‘ [^1] https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 for its high-quality embeddings across multiple languages, making it ideal for Italian health news [^2] After testing various chunk token sizes, the chunk length of 512 tokens with a 64-token overlap provide the best balance between retrieval performance and contextual continuity..
Figure 2: User-Agent Application Workflow. Text data is split into chunks for efficient retrieval, indexed using a Sentence Transformer model, and stored as dense vector embeddings in a vector store. The system retrieves relevant chunks based on query embeddings and passes them to an LLM for response generation.
For the retrieval part, I choose various LLMs handling the Italian language. LLaMA2 and LLaMA3 are generative models developed by Meta[49, 30], optimized through supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF). Mistral[18] is a 7-billion parameter model offering competitive performance in generating accurate responses. GPT-4[1], the latest multimodal model from OpenAI. Finally, Gemma and Gemma 2[46, 47], lightweight models from Google, are fine-tuned for Italian question-answering and summarization tasks.
The workflow illustrated in Figure 2 starts with retrieving relevant information from a large corpus of chunk news by embedding the user query into a vector space using the sentence transformer. The query embedding is matched against stored dense vector embeddings, and the top-K most similar document chunks are retrieved from the vector database. Next, the generation component uses the retrieved document chunks as context to produce a response. The LLM receives both the original query and the retrieved context, generating a natural language response grounded in factual data, thus reducing hallucinations.
4.3 Prompt Engineering Template
I develop a question-answering prompting template that guides the model in providing contextually relevant answers based on the information available in the provided context [^3] The list of the final set of Italian questions is available upon request.. An English version is shown in Table 4. This template ensures that the chatbot remains anchored to the information in the provided context. If the answer to a query is not found within the context, the chatbot will respond with: I’m sorry, I can’t help you based on the information I have”.
Table 4: Prompt Engineering - Example (English translation)
Prompt You are a chatbot that loves helping people! Given the following context section, answer the question using only the provided context. If you are unsure and the answer is not explicitly written in the documentation, respond by saying, ’I’m sorry, I can’t help you based on the information I have.’
Question How can myocardial infarction be prevented?
Context [1] The diagnosis of myocardial infarction on-site allows for rapid transportation by equipped ambulance, uncomfortable but not risky, directly to the hemodynamic lab, enabling coronary reopening at least 30 minutes earlier than if the patient arrives at the emergency room with their own means at the reference hospital.” – responsible for myocardial infarction [2] early infarction and saving lives with a simple blood test.” According to a press release from the Mauriziano’s press office. ”In the last ten years,” the press release explains, ”the diagnosis of acute coronary syndromes has progressively shifted attention from the electrocardiogram (always and essential) to the qualitative and quantitative biochemical definition of myocardial damage. The new method provides a more specific, sensitive, and early definition of myocardial damage with undeniable advantages in response speed to the pathology, especially when it presents subtly and unexpectedly. In short, it recognizes infarction practically in real-time.” [3] ”Remembering that acute myocardial infarction (AMI) is a time-dependent pathology,” Mura concludes, ”it is essential to prevent and initiate therapy, either pharmacological or interventional in the hemodynamic room, within the timeframes established by national protocols.
System Based on the context, there is no precise information on the symptoms indicating myocardial infarction. However, it mentions a new hospital technique capable of recognizing infarction in real-time, which is crucial given the time-dependent nature of the pathology.
5 Results
5.1 RAG Performance Analysis
The results are presented in Figures 3-4, showing the models’ performances across three levels of questioning. According to [8], a ROUGE score below 0.4 is considered low, between 0.4 and 0.5 is moderate, and above 0.5 is good.
The Gemma model shows the weakest performance across all metrics. Its ROUGE-1 and ROUGE-2 F1 scores fall below 0.4, indicating significant difficulty in capturing relevant unigrams and bigrams. The ROUGE-L score is similarly weak, suggesting issues with structural coherence. Additionally, Gemma’s BLEU and METEOR scores remain low, showing poor n-gram precision and recall. With no EM, Gemma struggles to generate sequences that match the reference.
In contrast, Gemma2 emerges as the top performer across all metrics. Its ROUGE-1 and ROUGE-2 F1 scores exceed 0.5, indicating strong unigram and bigram relevance. The high ROUGE-L score points to solid structural coherence, while the BLEU and METEOR scores suggest excellent n-gram precision and recall. Gemma2 also achieves notable EM rates, ranging between 40% and 50%, highlighting its strength in generating exact matches, which is critical for tasks requiring precise language generation.
GPT-4 delivers balanced performance across all metrics but shows variability across question levels. It follows closely behind Gemma2 in ROUGE, BLEU, and METEOR scores. However, a drop occurs at mid and high question levels, suggesting GPT-4 faces challenges with complex queries. Despite this, its BERTSCORE exceeds 0.9, and it also performs strongly in EM, rivaling Gemma2 in generating exact matches.
Llama2, Llama3, and Mistral exhibit similar patterns. While their F1 scores for all ROUGE variants are relatively low, mainly due to weak precision, their recall scores approach those of Gemma2 and GPT-4. Mistral maintains consistent performance across all question levels, but the Llama models struggle more with more difficult questions.
Overall, Gemma2 emerges as the most performant model, strictly followed by GPT-4, which remains competitive despite a slight decline in performance on complex queries. Mistral and the Llama models perform modestly across ROUGE, BLEU, and METEOR, with Mistral being more stable across all levels. Gemma lags significantly behind, particularly in ROUGE and BLEU. BERTSCORE remains consistently high across all models, with Gemma2 and GPT-4 achieving BERTSCOREs above 0.9.
Figure 3: ROUGE Metrics by Question Levels
Figure 4: BLEU, METEOR and EXACT MATCH Metrics by Question Levels
5.2 NMISS Performance Evaluation
Results for BLEU and METEOR NMISS scores are shown in Tables 5-6. Findings report the scores against the novel metric for all the models across the question levels. In both cases, NMISS scores generally surpass their standard counterparts, indicating enhanced performance across different difficulty levels.
Table 5: BLEU versus BLEU NMISS scores
BLEU score BLEU NMISS score
Model LOW MID HIGH LOW MID HIGH
gpt4 0.650 0.468 0.473 0.653 0.469 0.484
llama2 0.122 0.072 0.067 0.151 0.098 0.090
llama3 0.152 0.208 0.125 0.170 0.259 0.173
mistral 0.272 0.263 0.191 0.291 0.291 0.225
gemma 0.001 0.000 0.047 0.001 0.011 0.049
gemma2 0.654 0.578 0.604 0.668 0.626 0.664
33footnotetext: Note: Average scores. Higher values are highlighted in bold.
For both metrics, models like GPT-4 show minimal differences between standard and NMISS scores, indicating high consistency in performance across both metrics. However, models like Llama2, Llama3 and Mistral exhibit more significant improvements in their BLEU NMISS scores, suggesting that NMISS better captures the models’ ability to provide additional contextual information, especially in more challenging tasks. Notably, Gemma2 shows strong performance overall, with NMISS scores slightly higher than standard ones, particularly at mid and high levels, highlighting its ability to maintain accuracy and detail in complex scenarios. The increase in NMISS scores when models struggle to answer could be attributed to how NMISS rewards the inclusion of additional relevant information or contextual details. When a model does not provide a concise answer, it often compensates by providing surrounding context, explanations, or related information. While not a perfect match to the expected answer, this additional content contributes meaningfully to the overall response.
Table 6: METEOR versus METEOR NMISS scores
METEOR score METEOR NMISS score
Model LOW MID HIGH LOW MID HIGH
gpt4 0.714 0.526 0.565 0.714 0.526 0.567
llama2 0.443 0.325 0.291 0.453 0.330 0.310
llama3 0.415 0.442 0.336 0.429 0.455 0.355
mistral 0.431 0.485 0.451 0.436 0.488 0.456
gemma 0.010 0.012 0.073 0.011 0.019 0.080
gemma2 0.718 0.675 0.724 0.723 0.682 0.731
33footnotetext: Note: Average scores. Higher values are highlighted in bold.
Tables 7-9 compares ROUGE and ROUGE NMISS scores across different models and difficulty levels. Similar to the trends observed with BLEU and METEOR, the ROUGE NMISS scores generally outperform the standard ROUGE metrics, especially at higher difficulty levels. This consistent pattern across different metrics suggests that NMISS captures additional relevant information, rewarding models for providing more detailed and contextually accurate responses.
Specifically, for most models, ROUGE NMISS scores surpass the standard ROUGE-1 and ROUGE-L scores across all difficulty levels. This is consistent with the behavior observed for BLEU and METEOR. Interestingly, the ROUGE-2 NMISS scores do not consistently outperform the standard ROUGE-2 scores across all models and levels. This deviation from the trend seen in ROUGE-1 and ROUGE-L might be attributed to the focus of ROUGE-2 on bigram overlaps, which are less sensitive to the additional contextual information.
Table 7: ROUGE-1 versus ROUGE-1 NMISS scores
ROUGE-1 versus ROUGE-1 NMISS
Level Model P R F1 NMISS_P NMISS_R NMISS_F1
LOW gpt4 0.721 0.740 0.718 0.742 0.740 0.728
LOW llama2 0.201 0.712 0.282 0.342 0.717 0.421
LOW llama3 0.312 0.557 0.338 0.462 0.564 0.433
LOW mistral 0.434 0.489 0.421 0.531 0.492 0.466
LOW gemma 0.028 0.018 0.014 0.043 0.018 0.021
LOW gemma2 0.793 0.727 0.738 0.858 0.731 0.758
MID gpt4 0.552 0.529 0.530 0.603 0.531 0.539
MID llama2 0.153 0.510 0.211 0.280 0.511 0.308
MID llama3 0.356 0.605 0.384 0.558 0.610 0.488
MID mistral 0.434 0.547 0.447 0.547 0.550 0.498
MID gemma 0.017 0.018 0.015 0.049 0.022 0.026
MID gemma2 0.727 0.698 0.688 0.877 0.701 0.735
HIGH gpt4 0.585 0.603 0.582 0.647 0.604 0.611
HIGH llama2 0.156 0.477 0.211 0.301 0.485 0.330
HIGH llama3 0.271 0.432 0.295 0.491 0.452 0.401
HIGH mistral 0.364 0.537 0.396 0.499 0.539 0.455
HIGH gemma 0.084 0.084 0.076 0.111 0.091 0.095
HIGH gemma2 0.706 0.755 0.715 0.880 0.758 0.790
Table 8: ROUGE-2 versus ROUGE-2 NMISS scores
ROUGE-2 versus ROUGE-2 NMISS
Level Model P R F1 NMISS_P NMISS_R NMISS_F1
LOW gpt4 0.665 0.682 0.669 0.665 0.682 0.669
LOW llama2 0.131 0.568 0.191 0.131 0.568 0.191
LOW llama3 0.218 0.449 0.243 0.223 0.450 0.245
LOW mistral 0.337 0.409 0.340 0.338 0.409 0.340
LOW gemma 0.020 0.005 0.008 0.020 0.005 0.008
LOW gemma2 0.730 0.687 0.692 0.735 0.688 0.693
MID gpt4 0.490 0.486 0.485 0.490 0.486 0.485
MID llama2 0.084 0.353 0.120 0.085 0.353 0.121
MID llama3 0.270 0.472 0.284 0.273 0.473 0.286
MID mistral 0.326 0.453 0.349 0.333 0.454 0.351
MID gemma 0.004 0.003 0.004 0.004 0.003 0.004
MID gemma2 0.662 0.637 0.632 0.665 0.638 0.633
HIGH gpt4 0.493 0.505 0.496 0.494 0.505 0.496
HIGH llama2 0.075 0.280 0.106 0.076 0.281 0.107
HIGH llama3 0.155 0.295 0.179 0.163 0.296 0.181
HIGH mistral 0.242 0.412 0.281 0.244 0.413 0.282
HIGH gemma 0.067 0.056 0.060 0.067 0.056 0.060
HIGH gemma2 0.635 0.664 0.641 0.639 0.664 0.642
Table 9: ROUGE-L versus ROUGE-L NMISS scores
ROUGE-L versus ROUGE-L NMISS
Level Model P R F1 NMISS_P NMISS_R NMISS_F1
LOW gpt4 0.709 0.722 0.705 0.719 0.722 0.707
LOW llama2 0.171 0.657 0.245 0.206 0.658 0.284
LOW llama3 0.274 0.513 0.301 0.337 0.514 0.338
LOW mistral 0.390 0.454 0.385 0.440 0.456 0.406
LOW gemma 0.023 0.016 0.012 0.033 0.016 0.018
LOW gemma2 0.779 0.719 0.729 0.805 0.721 0.735
MID gpt4 0.538 0.519 0.519 0.575 0.520 0.523
MID llama2 0.129 0.460 0.180 0.165 0.461 0.214
MID llama3 0.321 0.555 0.345 0.373 0.555 0.377
MID mistral 0.377 0.500 0.398 0.413 0.502 0.413
MID gemma 0.011 0.013 0.009 0.022 0.014 0.014
MID gemma2 0.702 0.673 0.665 0.743 0.673 0.677
HIGH gpt4 0.551 0.562 0.546 0.573 0.562 0.554
HIGH llama2 0.120 0.406 0.167 0.159 0.408 0.201
HIGH llama3 0.227 0.382 0.252 0.293 0.389 0.285
HIGH mistral 0.316 0.489 0.353 0.356 0.489 0.372
HIGH gemma 0.079 0.079 0.072 0.092 0.084 0.082
HIGH gemma2 0.677 0.718 0.684 0.707 0.718 0.698
To provide a broad perspective on the various scores, Figures 5-6 illustrate the percentage of instances where NMISS outperforms standard metrics for all questions. For ROUGE-1 and ROUGE-L, NMISS consistently performs better, reflecting its ability to capture more contextually rich and detailed responses. This further highlights NMISS’s effectiveness in evaluating model performance, particularly in greater complexity scenarios.
The performance of ROUGE NMISS for precision across all question levels is as follows: for GPT-4 ROUGE NMISS outperforms ROUGE-1 by approximately 26% at the low level, 30% at the mid level, and 40% at the high level. ROUGE-L shows a similar improvement at all levels, while ROUGE-2 exhibits no notable change. For Llama2 and Llama3, ROUGE NMISS shows remarkable improvement in ROUGE-1, with gains between 80% and 85% across both low and high levels. ROUGE-L demonstrates more modest increases, ranging from 40% to 60%. For Mistral, NMISS provides more modest improvements, ranging from 4% to 65% in ROUGE-1 and similar results for ROUGE-L. Gemma2 follows a similar pattern to GPT-4, with modest improvements across all levels. At the same time, Gemma shows little to no improvement with NMISS, indicating that its issues may stem from machine-related problems rather than contextual relevance.
For recall, similar trends are observed, though with different magnitudes. This is expected, as recall is more sensitive to the length of the context, but they remain noteworthy. GPT-4 gains 18% for ROUGE-1 at mid-level questions, Llama3 improves by 11% at low and 12% at high levels, and Mistral improves by 6% at the low level. Gemma shows no change at the low level but gains 6.1% at mid and high levels.
The same patterns for the models across the question levels apply to METEOR and BLEU percentage analysis, where the difference stands in magnitude. While NMISS heavily impacts BLEU, METEOR is less pronounced.
Figure 5: Percentage of Outperforming of ROUGE NMISS scores
Figure 6: Percentage of Outperforming of BLEU AND METEOR NMISS scores
Overall, the NMISS algorithm demonstrates the greatest improvement at the highest difficulty level, where the gap between NMISS and traditional metrics like BLEU and METEOR is most pronounced. This suggests that NMISS is particularly effective in advanced scenarios that require deeper comprehension and more complex information retrieval or when the model’s predictive power is limited. Notably, with GPT-4 or Gemma2 being the top performers, they benefit the least from NMISS, indicating that answers are already highly contextual and do not need additional elaboration. In contrast, middle-range models like Llama2, Llama3, and Mistral show the most significant gains, as NMISS helps to capture their ability to provide richer contextual information. Gemma sees little to no improvement, indicating that its issues are likely machine-related rather than due to contextual relevance.
These results suggest that when models struggle to provide exact answers, they compensate by offering more detailed contextual information. Importantly, this does not result in hallucinations but rather reflects the models’ ability to maintain accuracy and completeness in the surrounding context. In such cases, even peripheral contexts can contribute to higher NMISS scores, improving the overall response quality. Conversely, where questions are more straightforward, the gap between NMISS and traditional metrics narrows. This indicates that less contextual elaboration is required, and the models perform similarly across both metrics. In summary, NMISS tends to reflect the models’ ability to provide richer, more informative answers in complex situations, which is less evident at lower difficulty levels where responses tend to be more concise and direct, reducing the gap between NMISS and traditional metrics.
6 Conclusions
In this paper, I investigate hallucinations in LLMs using a combination of detection and mitigation techniques. I evaluate a user-agent chatbot interaction with Italian health news articles through a question-answering framework based on the RAG architecture. Models including Gemma, Gemma2 GPT-4, Llama2, Llama3, and Mistral are analyzed. To enhance traditional evaluation metrics like ROUGE, METEOR, and BLEU, I introduce the NMISS scoring system, which emphasizes contextual relevance in model-generated responses. NMISS allows the identification of instances where traditional metrics undervalue contextually rich answers that deviate from the reference.
Regarding performance, Gemma2 and GPT-4 consistently outperformed other models. Gemma2 demonstrated strong results across ROUGE, METEOR, and BLEU, making it a top choice for healthcare-related chatbot applications. GPT-4 excelled in EM accuracy, highlighting its ability to generate sequences closely aligned with reference answers.
The mid-range models Llama2, Llama3, and Mistral show similar performance patterns. Although their F1 scores across all ROUGE variants are lower due to weaker precision, their recall scores are close to those of Gemma2 and GPT-4, indicating their ability to capture a significant amount of relevant content despite generating less precise outputs. Mistral maintains consistent performance across all difficulty levels, while Llama2 and Llama3 struggled with more complex queries.
NMISS application reveals that when a model struggles to provide exact answers, it often compensates by delivering richer contextual information. This does not result in hallucinations but highlights the LLM’s capacity to provide valuable information embedded in the context, even when direct answers are challenging to generate. This demonstrates the importance of context as a carrier of valuable information. NMISS consistently outperforms traditional ROUGE-1 and ROUGE-L scores across all question levels, and similar trends are observed for BLEU and METEOR.
It is important to note that high-performing models like Gemma2 and GPT-4 benefit less from NMISS, which is expected. These models already provide highly precise and contextually accurate responses. In contrast, NMISS proves particularly beneficial for mid-tier models like Llama2, Llama3, and Mistral, which often generate contextually richer responses when faced with more challenging queries. These models benefit from NMISS because it captures and rewards their ability to provide additional relevant context, even if their precision is lower. On the other hand, Gemma shows slight improvement across all metrics, suggesting that its hallucination issues are machine-intrinsic and less related to context-awareness, which NMISS is designed to evaluate.
While RAG models mitigate hallucinations by grounding outputs in external data, NMISS refines the evaluation process. It helps distinguish “false positives” — cases where traditional metrics flag hallucinations due to deviations from the reference, even though the response is contextually accurate. NMISS reduces penalties for models that provide relevant contextual information, even when their answers do not match the reference. This distinction underlines the complementary nature of RAG and NMISS: RAG acts as a tool for preventing hallucinations, while NMISS improves how I evaluate context and relevance in model output, serving as a diagnostic tool for better understanding hallucinations.
Future research could explore extending this framework to more complex datasets, tasks, and languages, further advancing hallucination mitigation and improving LLM performance in real-world applications.
Acknowledgements
This publication was produced while attending the PhD program in Economics and Business at the University of Cagliari, Cycle 38, with the support of a scholarship financed by the Ministerial Decree no. 351 of 9th April 2022, based on the NRRP - funded by the European Union - NextGenerationEU - Mission 4 “Education and Research”, Component 1 “Enhancement of the offer of educational services: from nurseries to universities” - Investment 3.4 “Advanced teaching and university skills”.